A loop of wire is placed in a magnetic field of uniform strength

3.0 T that is parallel to the normal of the loop. The area of the loop begins

shrinking at the rate of 2.0 m^2/s. What is the magnitude of the emf

induced in the loop while it is shrinking?


1
Expert's answer
2022-05-15T17:03:16-0400
"V=\frac{\Delta\Phi}{\Delta t}=\frac{B\Delta A}{\Delta t}"

"V=3.0*2.0=6.0\:\rm V"